can you help me to find three numbers which are in arithmetic series and in geometric series and the common difference or common ratio should not be zero
Answer by stanbon(75887)

can you help me to find three numbers which are in arithmetic series and in geometric series and the common difference or common ratio should not be zero
------
Arithmetic Sequence:: a, a+d, a+2d
--------
If that is also a geometric sequence,
(a+d)/a = (a+2d)/(a+d)
-------
(a+d)^2 = a(a+2d)
------
a^2 + 2ad + d^2 = a^2 + 2ad
-----
d^2 = 0
d = 0
-----
Then the arithmetic sequence = a,a,a
which is not an aritmetic sequence.
------
Answer to your question. There is no answer.
Cheers,
Stan H.
------------------------